# MS-IMAP - A Multi-Scale Graph Embedding Approach for Interpretable Manifold Learning

### v4

@@ Abstract
Deriving meaningful representations from complex, high-dimensional data in unsupervised settings is crucial across diverse machine learning applications. This paper introduces a framework for multi-scale graph network embedding based on spectral graph wavelets that employs a contrastive learning approach. We theoretically show that in Paley-Wiener spaces on combinatorial graphs, the spectral graph wavelets operator provides greater flexibility and control over smoothness compared to the Laplacian operator, motivating our approach. An additional key advantage of the proposed embedding is its ability to establish a correspondence between the embedding and input feature spaces, enabling the derivation of feature importance. We validate the effectiveness of our graph embedding framework on multiple public datasets across various downstream tasks, including clustering and unsupervised feature importance.
@@ 1 Introduction
Graph Embeddings and Manifold Learning [29, 3, 34, 35, 5] are pivotal in analyzing complex data structures prevalent in diverse machine learning applications. The emergent representations from these techniques offer insights into the underlying data structure in conjunction with downstream tasks like clustering and visualization and are particularly valuable when labels are unavailable or unreliable. However, limitations of these techniques include reliance on low frequencies of the graph Laplacian and $k$-Nearest Neighbors graph connectivity, which provide limited information, including in recent contrastive learning methods such as Uniform Manifold Approximation and Projection (UMAP). To address these limitations, this paper proposes three main contributions.
(1) A framework employing multi-scale graph representation using 3D based tensor and optimization using contrastive learning which enhances the expressiveness of an embedding by optimizing low and high-frequency information, capturing intricate data details. (2) An application of manifold interpretation to feature importance allows each dimension of our proposed embedding to be uniquely traced back and associated with the original features. (3) A characterization of the theoretical properties of the representation power of the spectral graph wavelets (SGW) operator by considering functions sampled from the Paley-Wiener spaces [27] on combinatorial graphs and a demonstration of the flexibility and better control over smoothness properties of the SGW operator in comparison to the Laplacian operator.
A key distinction between our approach and many manifold learning methods such as UMAP is that we leverage multi-scale graph representation (using SGW) to increase the dimensionality of the data thus enhancing data representation learning. We use stochastic gradient descent (SGD) to optimize the embedding space and employ a novel contrastive learning approach using a 3D-tensor to capture both low and high frequencies.
Generally, manifold learning methods perform nonlinear dimensionality reduction by embedding the data into lower-dimensional spaces, although some approaches may embed the data into higher dimensions [16, 31]. A common limitation of these and other graph embedding methods is their lack of direct connections to the input features, which restricts their ability to produce interpretable results. This can be crucial in fields like finance, where behavioral analysis must be summarized using key features to guide decisions, or in biology, where interpretable features are essential for understanding complex systems. Our proposed approach establishes a mapping between the original features and the constructed optimized embeddings and enables the measurement of the importance of the original features with respect to the embedding space. To the best of our knowledge, current manifold learning techniques produce embeddings lacking explicit interpretation. In contrast, our approach demonstrates the ability to generate interpretable embedding representations, while also being competitive with the state-of-the-art graph embeddings across applications such as finance, vision, and biology.
@@ 2 Related Work
Extensive efforts have been dedicated to exploring non-linear dimensionality reduction aimed at achieving a low-dimensional embedding that preserves the underlying manifold structure. Early manifold learning techniques include Locally Linear Embedding (LLE) [29], Laplacian Eigenmaps [3], Isomap [34], Diffusion Maps [4], and t-distributed Stochastic Neighbor Embedding (t-SNE) [35].
Balancing local and global structure in the resulting embeddings is challenging in the presence of noise. To address the challenge, researchers have introduced de-noising methods [20, 12, 8, 11, 7, 13, 9, 10]. These techniques seek to enhance algorithms robustness, thereby broadening the applicability of these techniques in real-world scenarios.
UMAP [25] has proven effective in providing vivid visualizations and cluster separation, yet its reliance on Laplacian Eigenmaps initialization and negative sampling optimization has limitations. Laplacian Eigenmaps initialization focuses on low-frequency signal overlooking higher-frequency patterns. It uniformly samples edges without considering topological importance, overemphasizing graph connectivity and introducing further distortion.
Recent work [6] further highlights that negative sampling has a significant impact on the effective loss and introduces distortions with t-SNE and UMAP which motivates the exploration of alternate sampling schemes. Other relevant studies include [38] which investigates competing negative sampling strategies, [32] on generalization error bounds , and [23] on topological and sampling bias due to high degree nodes.
The lack of explicit mapping linking the original high dimensional dataset to its low dimensional embedding has prevented their use in areas where interpretability is critical. Establishing this link clearly would lead to a better understanding of the original feature space and improve feature selection, understanding of the resulting visualizations, and robustness to noise [21, 15, 39].
@@
SGW serve as an efficient tool for localization in both the spectral and vertex domains, playing a crucial role in our method (see Section 3.1 for details). SGW are constructed by applying a set of bandpass filters to the eigenvalues of the graph Laplacian, enabling multi-resolution representation of graph signals. Also related to our approach are graph embedding and multi-scale framework using Diffusion Wavelets [28] that embed the data in higher dimensions such as the Geometric Scattering for Graph Data Analysis [17] and the Graph Scattering Transform [16] that extract graph features, which are then employed in graph classification tasks.
@@ 2.1 Preliminaries
Consider a set of points $\mathbf{x}=\left\{\mathbf{x}_{i}\right\},\,i=1,...N,\mathbf{x}_{i}\in\mathbb{R}^{D}$, which are sampled from an unknown manifold $M\subset\mathbb{R}^{D}$. An undirected, weighted graph $G=(V,\mathbf{W})$, is constructed over $\mathbf{x}$, where $V$ corresponds to the nodes and $\mathbf{W}$ to the set of edges on the graph. The adjacency matrix $\mathbf{W}=(w_{ij})$ consists of the weights $w_{ij}$ between node $i$ and node $j$. The weights can be chosen using different techniques, such as the Gaussian kernel function between the points $\mathbf{x}_{i},\mathbf{x}_{j}$ using the $k$ nearest neighbors of $\mathbf{x}_{i}$, kNN$(\mathbf{x}_{i})$ (which denotes the $k$ nearest neighbors) or adaptive graph construction as proposed in [25]. $\mathbf{L}$ denotes the combinatorial graph Laplacian, defined as $\mathbf{L}=\mathbf{D}-\mathbf{W}$, with $\mathbf{D}$ the diagonal degree matrix with entries $d_{ii}=d(i)$. The degree $d(i)$ of vertex $i$ is defined as the sum of weights of edges that are connected to $i$. The normalized Laplacian is defined as ${\mathbf{L}}_{N}=\mathbf{D}^{-\frac{1}{2}}\mathbf{L}\mathbf{D}^{-\frac{1}{2}}=\mathbf{I}-\mathbf{D}^{-\frac{1}{2}}\mathbf{W}\mathbf{D}^{-\frac{1}{2}}$ and its real eigenvalues are in the interval $[0,2]$. The eigenvalues and eigenvectors of $\mathbf{L}$ are $\lambda_{1},\ldots,\lambda_{N}$ and $\mathbf{\phi}_{1},\ldots,\phi_{N}$, respectively. In this work we use the normalized Laplacian $\mathbf{L}_{N}$, which from now on we refer to as $\mathbf{L}$ for simplicity. We will denote the matrix representations of the eigenvectors and eigenvalues of the graph Lapacian $\mathbf{L}$ as $\Phi,\lambda$, respectively. Our approach assumes that all graph signals $f_{r}\in\mathbb{R}^{N}$ used to create the embedding correspond to the input coordinates, $\mathbf{x}=\left(f_{1},f_{2},..,f_{r},..f_{D}\right)$. The output embedding space provides a multi-scale representation approximating the intrinsic manifold coordinates.
@@ 2.2 Multi-scale representations using SGW
In the last two decades, several multi-scale representations using irregular graphs have been suggested, including SGW [18] and Diffusion Wavelets [28]. In this work we focus on the multi-scale graph transform utilizing SGW. SGW provides a natural way of trading off between spectral and spatial resolution, since the SGW coefficients are localized in both domains. These wavelets are constructed using a kernel function $g(\mathbf{L})$ operator, which acts on a function by modulating each of its Fourier modes [18], that helps in capturing a trade off between vertex (spatial) and spectral localization. Spatial localization in the graph domain is implicitly controlled by a single wavelet scale parameter defined in the spectral domain, such that the more the vertex is localized in the vertex domain, the spectral band is wider. To represent a signal ${f}\in\mathbb{R}^{N}$ in multiple scales $S=[s_{1},s_{2},..s_{K}]$, the SGW transform is constructed as follows: Assume that $g(\lambda)$ is a filter in the spectral domain. Let $\delta_{i}\in\mathbb{R}^{N}$ be the delta function centred at the vertex $i\in G:\delta_{i}(j)=1\,\,\mbox{if}\,\,i=j$, and $\delta_{i}(j)=0$ otherwise. Given a bandpass filter $g(\lambda)$ and a wavelet centered on node $i$ at scale $s$, the wavelets $\psi(s,i),i=1,..N$ are calculated by applying them to the delta function on a single vertex $i$, given by $\psi(s,i)=\Phi g(s\lambda)\Phi^{T}\delta_{i}$. The value of $\psi(s,i)$ with respect to a vertex $m$ can be written as $\psi(s,i)(m)=\sum_{l=1}^{N}g(s\lambda_{l})\phi_{l}^{*}(i)\phi_{l}(m)$. Given a graph signal ${f}\in\mathbb{R}^{N}$, the SGW coefficient at node $i$ and scale $s$ can be expressed as follows:
@@
Fast computation using Chebyshev polynomials: Directly computing the SGW coefficients is computationally intensive - $O(N^{3})$ for $N$ points. Instead, Hammond et al. [18] suggested computing the SGW using a fast algorithm based on approximating the scaled generating kernels through low-order polynomials of $\mathbf{L}$ applied to the input data (details in supplementary material).
@@ 3 Our Proposed framework: Multi-Scale IMAP
In this section, we introduce Multi-scale IMAP, a framework for interpretable embedding via manifold learning utilizing a multi-scale graph representation. This approach enables us to maintain global regularity and preserve local structure without sacrificing interpretability. Our method imposes a differentiable structure on the mapping $h:\mathbf{x}\rightarrow\psi_{\mathbf{x}}$, supported on a discrete graph $G=(V,\mathbf{W})$, where $\psi_{\mathbf{x}}$ represents the encoded multi-scale graph transform of $\mathbf{x}$. Multi-scale IMAP consists of two main components:
Step 1: Features representation encoding using multi-scale graph representation.
In this step, our approach constructs a multi-scale graph representation by incorporating feature signals through the SGW Transform across multiple scales and graph frequencies. We suggest a 3D tensor-based approach (detailed in Algorithm 1), where we construct a 3D tensor with dimensions ${K}\times{N}\times{D}$ to provide an effective representation for encoding multi-scale transform. We leverage the encoded 3D tensor structure to align the optimized embedding with all scales simultaneously to enforce differentiable structure of the transformed features. Note that an alternative would be to use simple concatenation, all of the features transformed for all scales are represented by concatenating all the SGW coefficients for each point in a vector form. For completeness we also describe this alternative in the Appendix. The 3D tensor-based approach (detailed in Algorithm 1) provides the most effective representation for the following reasons:
@@
(i) Joint and simultaneous optimization: using the 3D tensor-based method enables the joint and simultaneous optimization of all scales for each sample by leveraging the SGW coefficients of each spectral band within the cross-entropy loss framework while adhering to the fixed graph structure (Equation (3.2)). This approach effectively balances coarse and fine-grain resolutions, with summation performed only after the contrastive learning optimization process is complete. By smoothing both low- and high-frequency information jointly and simultaneously with respect to the graph structure, this method reduces the reliance on threshold-based techniques for removing high-frequency spectral bands, which may be less efficient. Jointly optimizing across scales ensures consistency in the optimization process for all scales with respect to a fixed pair of nodes. Although there may be correlations between overlapping SGWs corresponding to different spectral bands, each scale is associated with a distinct bandwidth of frequencies, effectively capturing graph structures at varying levels of neighborhood information. By integrating multiple scales, SGWs capture neighborhood information continuously across varying granularities, enabling a representation that strikes a balance between local and global structural details.
@@
(ii) Enhanced Interpretability: the proposed representation establishes a one-to-one correspondence between input features and the final embedding features, making feature interpretation more straightforward and intuitive. This is demonstrated in Section 6, where we show that the embeddings produced enable effective feature importance rankings. By ensuring this direct correspondence, our approach facilitates meaningful interpretation of the embedding space. This combination of optimization efficiency and enhanced interpretability highlights the clear advantages of the 3D tensor-based representation.
Step 2: Optimization design using multi-scale network structure.
We integrate the multi-scale representation within the SGD optimization framework, simultaneously leveraging both low and high-frequency information. This integration leads to fine-grained manifold regularization and improved robustness in downstream tasks.

### v5

@@ Abstract
Deriving meaningful representations from complex, high-dimensional data in unsupervised settings is crucial across diverse machine learning applications. This paper introduces a framework for multi-scale graph network embedding based on spectral graph wavelets that employs a contrastive learning approach. We theoretically show that in Paley-Wiener spaces on combinatorial graphs, the spectral graph wavelets operator provides greater flexibility and control over smoothness compared to the Laplacian operator, motivating our approach. A key advantage of the proposed embedding is its ability to establish a correspondence between the embedding and input feature spaces, enabling the derivation of feature importance. We validate the effectiveness of our graph embedding framework on multiple public datasets across various downstream tasks, including clustering and unsupervised feature importance.
@@ 1 Introduction
Graph Embeddings and Manifold Learning ([29, 3, 34, 35, 5]) play a pivotal role in analyzing complex data structures encountered in a wide range of machine learning applications. The representations learned by these techniques uncover the intrinsic geometry of the data and support downstream tasks such as clustering and visualization—especially in scenarios where labels are unavailable or unreliable. Most manifold learning methods typically perform nonlinear dimensionality reduction to embed data into lower-dimensional spaces, while some also explore high-dimensional embeddings ([16, 31]). Nonetheless, most such methods lack a direct connection to the input features, preventing the evaluation of individual feature contributions and resulting in embeddings with limited interpretability. This limitation is critical in domains like finance, where key behavioral indicators must be traced back to individual features, or in biology, where interpretability is vital for uncovering mechanisms within complex systems. To address these limitations, this paper makes three primary contributions:
1. We introduce a framework that leverages multi-scale graph representation using a 3D tensor and contrastive learning. This enhances the expressiveness of the embedding by jointly optimizing over both low- and high-frequency components, enabling the capture of fine-grained data structure.
2. We propose a manifold-based interpretation method for feature importance, allowing each dimension of the learned embedding to be directly linked back to a unique original feature.
3. We characterize the theoretical representational capacity of the Spectral Graph Wavelet (SGW) operator by analyzing functions from Paley-Wiener spaces ([27]) on combinatorial graphs. This analysis demonstrates that SGW provides greater flexibility and smoother control compared to traditional Laplacian-based operators.
Our method, MS-IMAP, is designed to associate each embedding coordinate with a unique input feature, thereby enabling the estimation of feature importance relative to the learned embedding space. This interpretability arises from a deliberate design choice: each embedding coordinate corresponds to a single, unique input feature and is optimized through spectral graph wavelets to align with the underlying graph structure. Unlike standard manifold learning and graph embedding methods, where each embedding dimension typically blends information from all input features, MS-IMAP preserves the semantic identity of each feature by assigning it to a distinct embedding dimension.
Another key distinction of our approach—compared to both classical ([3]) and modern techniques ([25]), which typically rely on the low-frequency components of the graph Laplacian and thus capture only coarse structural information—is the use of multi-scale graph representations via Spectral Graph Wavelets (SGW) within a contrastive learning framework. This allows our method to retain or even expand the dimensionality of the data, yielding richer and more expressive representations while preserving interpretability. We further employ stochastic gradient descent (SGD) with a novel 3D-tensor-based contrastive objective to optimize embeddings that capture both global and local graph structure. We empirically demonstrate that applying feature importance techniques to MS-IMAP embeddings yields more informative feature subsets for clustering than applying them to raw input data. This provides strong evidence of semantic alignment—often absent in nonlinear methods such as deep networks or conventional manifold learning. MS-IMAP achieves interpretable and semantically faithful embeddings, while remaining competitive with state-of-the-art graph embedding techniques across a range of tasks.
@@ 2 Related Work
Extensive efforts have been dedicated to exploring manifold learning methods that aim to produce low-dimensional embeddings while preserving the intrinsic structure of high-dimensional data. Classical manifold learning techniques include Locally Linear Embedding (LLE) [29], Laplacian Eigenmaps ([3]), Isomap ([34]), Diffusion Maps ([4]), and t-distributed Stochastic Neighbor Embedding (t-SNE) ([35]).
One of the central challenges in these methods is maintaining a balance between local and global structure, especially in the presence of noise. To address this, various denoising strategies have been developed ([20, 12, 8, 11, 7, 13, 9, 10]), aiming to improve robustness and extend the applicability of these techniques to real-world, noisy datasets.
UMAP ([25]) has shown strong empirical performance for visualization and cluster separation. It typically employs Laplacian Eigenmaps initialization, which emphasizes low-frequency graph signals while neglecting higher-frequency patterns that may carry important information.
Recent work ([6]) has shown that negative and positive sampling in UMAP significantly influences the effective loss function. These insights have led to studies on alternative sampling schemes ([38]), bounds on generalization error in embedding methods ([32]), and biases introduced by high-degree nodes during graph construction and sampling ([23]).
Manifold learning and non-linear dimensionality reduction serve purposes beyond visualization; one of their key goals is to learn representations that capture intrinsic geometric structure and approximate geodesic distances. The learned embedding space dimensionality can go beyond two or three dimensions. However, increasing dimensionality does not necessarily improve robustness: noise and the curse of dimensionality can introduce instability. These methods also suffer from computational limitations. Moreover, to compute embedding in higher dimensions many algorithms—such as Laplacian Eigenmaps and UMAP—require computing large portions of the Laplacian spectrum, making them computationally intensive for large datasets. Our method addresses these issues by producing interpretable and efficient embeddings while leveraging spectral information across multiple frequency scales.
A key limitation of many manifold learning and graph embedding methods is the absence of an explicit and interpretable mapping, which prevents evaluating the contribution of each input feature to the resulting low-dimensional embedding. This gap restricts their utility in domains where interpretability is crucial. Providing such a link can enable improved feature selection, clearer interpretation of embeddings, and greater robustness to noise ([21, 15, 39]).
@@
Spectral Graph Wavelets (SGW) are central to our method (see Section 3.1), offering localization in both spectral and vertex domains. SGW apply a family of bandpass filters to the graph Laplacian’s spectrum, allowing for multi-resolution analysis of graph signals and providing richer representations across frequency bands. Our approach is also related to graph embedding methods that exploit multiscale structure, such as Diffusion Wavelets ([28]), Geometric Scattering for Graph Data Analysis ([17]), and the Graph Scattering Transform ([16]), which extract graph features for downstream tasks like classification. These methods often embed data in higher-dimensional spaces to improve representation power.
@@ 2.1 Preliminaries
Consider a set of points $\mathbf{x}=\left\{\mathbf{x}_{i}\right\},i=1,...N,\mathbf{x}_{i}\in\mathbb{R}^{D}$, where each $\mathbf{x}_{i}\in\mathbb{R}^{D}$, sampled from an unknown smooth manifold $M\subset\mathbb{R}^{D}$. To model the geometric structure of the data, we construct an undirected, weighted graph $G=(V,\mathbf{W})$ over the point set $\mathbf{x}$, where $V$ is the set of nodes and $\mathbf{W}=(w_{ij})$ is the weighted adjacency matrix encoding edge weights between nodes. The weights $w_{ij}$ can be computed using various techniques. A common approach is to apply a Gaussian kernel to pairs of points $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ within the $k$ nearest neighbors of $\mathbf{x}_{i}$, denoted as $\text{kNN}(\mathbf{x}_{i})$. In this work, we adopt the adaptive Gaussian kernel construction proposed in [25], though other graph construction methods may also be employed. Let $\mathbf{L}$ denote the combinatorial graph Laplacian defined as $\mathbf{L}=\mathbf{D}-\mathbf{W}$, where $\mathbf{D}$ is the diagonal degree matrix with entries $d_{ii}=d(i)$, and the degree $d(i)$ of vertex $i$ is the sum of the weights of all edges incident to $i$. We use the normalized graph Laplacian, defined as:
$\mathbf{L}_{N}=\mathbf{D}^{-1/2}\mathbf{L}\mathbf{D}^{-1/2}=\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{W}\mathbf{D}^{-1/2}$
which has real eigenvalues in the interval $[0,2]$. For simplicity, we refer to $\mathbf{L}_{N}$ as $\mathbf{L}$ for the remainder of the paper. Let $\lambda_{1},\ldots,\lambda_{N}$ and $\boldsymbol{\phi}_{1},\ldots,\boldsymbol{\phi}_{N}$ denote the eigenvalues and corresponding eigenvectors of $\mathbf{L}$. We denote the matrix of eigenvectors as $\Phi\in\mathbb{R}^{N\times N}$ and the diagonal matrix of eigenvalues as $\Lambda=\text{diag}(\lambda_{1},\ldots,\lambda_{N})$. In our setting, each input coordinate $f_{r}\in\mathbb{R}^{N}$ — viewed as a graph signal—is defined across the $N$ nodes of the graph. The original data matrix $\mathbf{x}$ can thus be expressed as a collection of such signals: $\mathbf{x}=\left(f_{1},f_{2},..,f_{r},..f_{D}\right)$. Our goal is to derive a multi-scale embedding from these signals that approximates the intrinsic manifold coordinates, using spectral information derived from the graph Laplacian.
@@ 2.2 Multi-scale representations using SGW
In the past two decades, several multi-scale representations for data on irregular graphs have been developed, including Spectral Graph Wavelets (SGW) [18] and Diffusion Wavelets [28]. In this work, we focus on the SGW-based multi-scale graph transform. SGW offers a principled way to balance spectral and spatial resolution, as its coefficients are localized in both domains. These wavelets are constructed using a kernel function operator $g(\mathbf{L})$, which acts on a signal by modulating each of its spectral components (i.e., Fourier modes) [18]. This design enables a trade-off between vertex-domain (spatial) and frequency-domain (spectral) localization. The spatial localization is implicitly controlled by a single scale parameter defined in the spectral domain: greater localization in the vertex domain corresponds to a broader spectral bandwidth. To represent a signal $f\in\mathbb{R}^{N}$ at multiple scales $S=[s_{1},s_{2},\ldots,s_{K}]$, the SGW transform is defined as follows. Let $g(\lambda)$ be a bandpass filter in the spectral domain. Let $\delta_{i}\in\mathbb{R}^{N}$ denote the delta function centered at vertex $i\in G$, where $\delta_{i}(j)=1$ if $i=j$, and $\delta_{i}(j)=0$ otherwise. A wavelet centered at node $i$ and scale $s$ is then given by:
$\psi(s,i)=\Phi g(s\lambda)\Phi^{T}\delta_{i}$
.The value of $\psi(s,i)$ at vertex $m$ can be written explicitly as:
$\psi(s,i)(m)=\sum_{l=1}^{N}g(s\lambda_{l})\phi_{l}^{*}(i)\phi_{l}(m)$
. Given a graph signal $f\in\mathbb{R}^{N}$, the SGW coefficient at node $i$ and scale $s$ is defined as:
@@
Fast computation using Chebyshev polynomials:
Direct computation of SGW coefficients is computationally expensive, requiring $O(N^{3})$ operations for $N$ nodes. To address this, Hammond et al. [18] proposed an efficient algorithm based on approximating the scaled generating kernels $g(s\lambda)$ using low-order Chebyshev polynomials of the Laplacian $\mathbf{L}$, which are applied directly to the input signal. This significantly reduces computational cost (see Appendix for further details).
@@ 3 Our Proposed framework: Multi-Scale IMAP
In this section, we introduce Multi-scale IMAP, a framework for interpretable embedding via manifold learning utilizing a multi-scale graph representation. This approach enables us to maintain global regularity and preserve local structure without sacrificing interpretability. Our method imposes a differentiable structure on the mapping $h:\mathbf{x}\rightarrow\psi_{\mathbf{x}}$, supported on a discrete graph $G=(V,\mathbf{W})$, where $\psi_{\mathbf{x}}$ represents the encoded multi-scale graph transform of $\mathbf{x}$. Multi-scale IMAP consists of two main steps: In Step 1, our approach constructs a multi-scale graph representation by incorporating feature signals through the SGW Transform across multiple scales and graph frequencies. We suggest a 3D tensor-based approach (detailed in Algorithm 1), where we construct a 3D tensor with dimensions ${K}\times{N}\times{D}$ to provide an effective representation for encoding multi-scale transform. We leverage the encoded 3D tensor structure to align the optimized embedding with all scales simultaneously to enforce differentiable structure of the transformed features. The 3D tensor-based approach (detailed in Algorithm 1) provides the most effective representation for the following reasons:
@@
(i) Joint and simultaneous optimization: using the 3D tensor-based method enables the joint and simultaneous optimization of all scales for each sample by leveraging the SGW coefficients of each spectral band within the cross-entropy loss framework while adhering to the fixed graph structure. This approach effectively balances coarse and fine-grain resolutions, with summation performed only after the contrastive learning optimization process is complete. By smoothing both low- and high-frequency information jointly and simultaneously with respect to the graph structure, this method reduces the reliance on threshold-based techniques for removing high-frequency spectral bands, which may be less efficient. Jointly optimizing across scales ensures consistency in the optimization process for all scales with respect to a fixed pair of nodes. Although there may be correlations between overlapping SGWs corresponding to different spectral bands, each scale is associated with a distinct bandwidth of frequencies, effectively capturing graph structures at varying levels of neighborhood information. By integrating multiple scales, SGWs capture neighborhood information continuously across varying granularities, enabling a representation that strikes a balance between local and global structural details.
@@
(ii) Enhanced Interpretability: the proposed representation establishes a one-to-one correspondence between input features and the final embedding features, making feature interpretation more straightforward and intuitive. This is demonstrated in Section 4, where we show that the embeddings produced enable effective feature importance rankings. By ensuring this direct correspondence, our approach facilitates meaningful interpretation of the embedding space. This combination of optimization efficiency and enhanced interpretability highlights the clear advantages of the 3D tensor-based representation. We note that an alternative to the 3D tensor-based representation approach is to use simple concatenation, where all features are transformed using SGW and each data point is represented by a vector formed by concatenating all corresponding SGW coefficients. For completeness we describe this alternative in the Appendix. In Step 2, We integrate the multi-scale representation within the SGD optimization framework, simultaneously leveraging both low and high-frequency information. This integration leads to fine-grained manifold regularization and improved robustness in downstream tasks.